Find the complete integral of the differential equation

p2- y3q = x2 -y2.


1
Expert's answer
2021-10-27T10:23:57-0400

"\frac{dx}{-f_p}=\frac{dy}{-f_q}=\frac{dz}{-pf_p-qf_q}=\frac{dp}{f_x+pf_z}=\frac{dq}{f_y+qf_z}"


"\frac{dx}{-2p}=\frac{dy}{y^3}=\frac{dz}{2p^2-y^3q}=\frac{dp}{-2x}=\frac{dq}{2y}"


"2xdx=2pdp"

"p^2=x^2+c"

"y=c_1"


"2ydy\/y^3=dq"

"q=-2\/y+c'"


"\frac{d(\sqrt{x^2+c})}{-2x}=\frac{d(-2/y+c')}{2y}"


case 1:


"-\frac{dx}{2\sqrt{x^2+c}}=\frac{dy}{y^3}"


"-1/2y^2=-ln(\sqrt{x^2+c}+x)/2+c'_2"


"c_2=y^2-ln(\sqrt{x^2+c}+x)"


"F(c_1,c_2)=F(y,y^2-ln(\sqrt{x^2+c}+x))=0"


case 2:


"\frac{-d(\sqrt{x^2+c})}{-2x}=\frac{d(-2/y+c')}{2y}"


"-1/2y^2=ln(\sqrt{x^2+c}+x)/2+c'_2"


"c_2=y^2+ln(\sqrt{x^2+c}+x)"


"F(c_1,c_2)=F(y,y^2+ln(\sqrt{x^2+c}+x))=0"